A sample of an ideal gas has a volume of 3.60 L at 13.80 °C and 1.80 atm. What is the volume of the gas at 19.00 °C and 0.991 atm?
Answer #1

we have:

P1 = 1.80 atm

P2 = 0.991 atm

V1 = 3.60 L

T1 = 13.8 oC

= (13.8+273) K

= 286.8 K

T2 = 19.0 oC

= (19.0+273) K

= 292 K

we have below equation to be used:

(P1*V1)/(T1) = (P2*V2)/(T2)

(1.8 atm*3.6 L)/(286.8 K) = (0.991 atm*V2)/(292.0 K)

V2 = 6.66 L

Answer: 6.66 L

Feel free to comment below if you have any doubts or if this answer do not work